# Context Volume Drives Performance: Tackling Domain Shift in Extremely Low-Resource Translation via RAG

### v1

@@ Abstract
Neural Machine Translation (NMT) models for low-resource languages suffer significant performance degradation under domain shift. We quantify this challenge using Dhao, an indigenous language of Eastern Indonesia with no digital footprint beyond the New Testament (NT). When applied to the unseen Old Testament (OT), a standard NMT model fine-tuned on the NT drops from an in-domain score of 36.17 chrF++ to 27.11 chrF++. To recover this loss, we introduce a hybrid framework where a fine-tuned NMT model generates an initial draft, which is then refined by a Large Language Model (LLM) using Retrieval-Augmented Generation (RAG). The final system achieves 35.21 chrF++ (+8.10 recovery), effectively matching the original in-domain quality. Our analysis reveals that this performance is driven primarily by the number of retrieved examples rather than the choice of retrieval algorithm. Qualitative analysis confirms the LLM acts as a robust “safety net,” repairing severe failures in zero-shot domains.
@@
David Samuel Setiawan, Raphaël Merx, Jey Han Lau The University of Melbourne {david.setiawan, raphael.merx}@student.unimelb.edu.au jeyhan.lau@unimelb.edu.au
@@ 1 Introduction
We systematically compare sentence-level retrieval (whole-sentence similarity) against word-level retrieval (aggregated source word matches) to assess their impact on the proposed hybrid pipeline. Our results indicate that while increasing context volume consistently yields gains across all strategies, the specific choice of retrieval algorithm is secondary. The final optimized system restores character-level fluency (chrF++) to in-domain levels, with remaining lexical gaps (spBLEU) largely attributable to stylistic divergence between the Old and New Testaments. Based on these findings, we summarize our primary contributions as follows:
@@ 2 Related Work
Standard NMT models are highly lexicalized, making them brittle when applied to distributions differing from their training data (Koehn and Knowles, 2017; Hu et al., 2019; Haddow et al., 2022). Recent studies quantify this severity even when moving from religious texts to general domain dictionary sentences, finding that models trained solely on scripture fail to generalize without target-domain data (Rowe et al., 2025). Our work addresses an even stricter constraint: an internal domain shift (NT-to-OT) in a language with no digital footprint, where no such target-domain data exists.
@@ 3.2 Models
  • NMT (Drafting): We use NLLB-200-distilled-600M (Costa-jussà et al., 2024). Despite its smaller size, its massive multilingual pre-training provides a robust initialization for fine-tuning on the limited Dhao NT data.
  • LLM (Post-Editing): We utilize Gemini 2.5 Flash for the post-editing stage. This model was selected for its large context window (enabling the ingestion of extensive RAG examples) and its cost-effectiveness for iterative experimentation.
@@ 4.1 The Hybrid NMT-LLM Pipeline
We fine-tune NLLB-200-distilled-600M (Costa-jussà et al., 2024) on the in-domain (NT) corpus to generate an initial hypothesis $y_{nmt}$, adopting the optimal hyperparameters from the eBible benchmark (Akerman et al., 2023), detailed in Appendix D.
@@
We post-edit the translation with Gemini 2.5 Flash (Comanici et al., 2025). The LLM receives a structured prompt containing:
  1. 1.
    The original source sentence $x$.
  2. 2.
    The NMT draft $y_{nmt}$.
  3. 3.
    A set of retrieved parallel sentences (sourced from both the NT and the grammar book).
  4. 4.
    A set of retrieved lexicon entries formatted as direct mappings (e.g., English Word (POS) $\rightarrow$ Dhao Word).
The composition of these retrieved contexts depends on the experimental setup. We evaluate parallel sentence retrieval and lexicon retrieval independently, but combine them in the final optimized system (see Section 5.3). The full prompt structure and integration details are provided in Appendix C. The model is instructed to compare the draft against the source and correct specific errors (e.g., hallucinations, repetition loops) only when necessary, rather than re-translating from scratch.
@@ 4.2 Baselines
    LLM Direct Translation: Gemini 2.5 translating directly from English to Dhao.
    • 0-shot: No context provided.
    • 5-shot: Context consists of 5 fixed, randomly selected sentences from the in-domain NT corpus. This serves as a control for "blind" few-shot learning.
@@
    LLM Post-Editing (No RAG): The hybrid pipeline without retrieval, relying solely on the LLM’s internal knowledge to correct the NMT draft.
    • 0-shot: Instructions only.
    • 5-shot: Context consists of the same 5 fixed, randomly selected NT sentences as the direct translation baseline.
@@ 5.3 Final System Performance
Interestingly, while the combined approach yields the highest spBLEU (19.88), the chrF++ score is slightly lower than the sentence-only configuration (35.28). This suggests a nuanced trade-off: the lexicon improves exact lexical matching (boosting BLEU), but the sheer volume of combined context may introduce slight syntactic noise that affects character-level fluency. Nevertheless, the combined model remains the most robust overall system.
@@ 5.5 Recommendations
Standard practices in low-resource NMT often involve randomly splitting available corpora (e.g., the Bible) into training and test sets (Vázquez et al., 2021; Marashian et al., 2025). While many papers assume that the Bible belongs to a single, religious domain, our analysis shows a marked domain shift within this text, demonstrating that measuring out-of-domain performance is possible even when only the Bible is available as parallel corpus. Therefore, we recommend that low-resource researchers take this into consideration instead of using all verses of the Bible for both train and test, opting instead for document-level holdouts (e.g., distinct books or Testaments) to avoid inflated performance estimates (Khiu et al., 2024).
@@ 6 Conclusion
Crucially, we find that ontext volume, rather than retrieval algorithm, is the primary driver of performance. We observe that distinct retrieval strategies (lexical vs. semantic) converge to comparable quality levels when normalized for volume. By validating these trade-offs on a language with no digital footprint, we provide a scalable blueprint for accelerating the translation of the Old Testament for thousands of low-resource languages worldwide.
@@ 7 Limitations
This work utilizes data derived from the Dhao Grammar book and the Dhao Bible translation (New Testament). These resources were accessed from publicly available repositories and are used strictly for non-commercial research purposes to advance NLP for low-resource languages. We acknowledge the intellectual property of the original authors and translators. The datasets constructed in this work are released with a license restricted to academic and research use to prevent commercial exploitation of indigenous cultural resources.

### v2

@@ Abstract
Neural Machine Translation (NMT) models for low-resource languages suffer significant performance degradation under domain shift. We quantify this challenge using Dhao, an indigenous language of Eastern Indonesia with no digital footprint beyond the New Testament (NT). When applied to the unseen Old Testament (OT), which exhibits a 3x increase in OOV rate (25.9%) and distinct thematic divergence from the New Testament (NT) training data, a standard NMT model fine-tuned on the NT drops from an in-domain score of 36.17 chrF++ to 27.11 chrF++. To recover this loss, we introduce a hybrid framework where a fine-tuned NMT model generates an initial draft, which is then refined by a Large Language Model (LLM) using Retrieval-Augmented Generation (RAG). The final system achieves 35.21 chrF++ (+8.10 recovery), effectively matching the original in-domain quality. Our analysis reveals that this performance is driven primarily by the number of retrieved examples rather than the choice of retrieval algorithm. Qualitative analysis confirms the LLM acts as a robust “safety net,” repairing severe failures in zero-shot domains.
@@
David Samuel Setiawan, Raphaël Merx, Jey Han Lau The University of Melbourne {david.setiawan, raphael.merx}@student.unimelb.edu.au laujh@unimelb.edu.au
@@ 1 Introduction
We systematically compare sentence-level retrieval (whole-sentence similarity) against word-level retrieval (aggregated source word matches) to assess their impact on the proposed hybrid pipeline. Our results indicate that while increasing context volume consistently yields gains across all strategies, the specific choice of retrieval algorithm is secondary. The final optimized system restores character-level overlap (chrF++) to in-domain levels, though a gap remains in subword-level similarity (spBLEU), likely due to the higher stylistic and lexical divergence of the Old Testament. Based on these findings, we summarize our primary contributions as follows:
@@ 2 Related Work
Standard NMT models are highly lexicalized, making them brittle when applied to distributions differing from their training data (Koehn and Knowles, 2017; Hu et al., 2019; Haddow et al., 2022). This brittleness often manifests as catastrophic hallucinations or fluent but unfaithful outputs when the model encounters out-of-domain data Müller et al. (2020); Raunak et al. (2021). Specifically, such shifts have been shown to exacerbate the model’s reliance on training-domain priors, leading it to ignore source constraints in favor of frequently observed sequences from the training set Wang and Sennrich (2020).
@@ 3.2 Models
  • NMT (Drafting): We use NLLB-200-distilled-600M (Costa-jussà et al., 2024). We specifically select this distilled version over larger variants (e.g., 1.3B or 3.3B) to prioritize faster inference speeds. This allows for rapid iteration during experimentation while still leveraging the model’s massive multilingual pre-training, which provides a robust initialization for fine-tuning on the limited Dhao NT data.
  • LLM (Post-Editing): We utilize Gemini 2.5 Flash (Comanici et al., 2025) for the post-editing stage. This model was selected for its large context window (enabling the ingestion of extensive RAG examples) and its cost-effectiveness for iterative experimentation.
@@ 4.1 The Hybrid NMT-LLM Pipeline
We fine-tune the NMT model described in Section 3.2 on the in-domain (NT) corpus to generate an initial hypothesis $y_{nmt}$. We adopt the optimal hyperparameters from the eBible benchmark (Akerman et al., 2023), as detailed in Appendix D.
@@
We post-edit the translation with Gemini 2.5 Flash (Comanici et al., 2025). The LLM receives a structured prompt containing: (1) the original source sentence $x$; (2) the NMT draft $y_{nmt}$; (3) a set of retrieved parallel sentences (sourced from both the NT and the grammar book); and (4) a set of retrieved lexicon entries formatted as direct mappings (e.g., English Word (POS) $\rightarrow$ Dhao Word). The composition of these retrieved contexts depends on the experimental setup. We evaluate parallel sentence retrieval and lexicon retrieval independently, but combine them in the final optimized system (see Section 5.3). The full prompt structure and integration details are provided in Appendix C. The model is instructed to compare the draft against the source and selectively correct NMT failures like hallucinations or repetition loops only when necessary rather than re-translating from scratch.
@@ 4.2 Baselines
    LLM Direct Translation: Gemini 2.5 translating directly from English to Dhao in 0-shot (no context) and 5-shot (5 fixed, randomly selected NT sentences) settings
@@
    LLM Post-Editing (No RAG): The hybrid pipeline without retrieval, relying on internal LLM knowledge to correct the NMT draft in 0-shot and 5-shot settings.
@@ 5.3 Final System Performance
Interestingly, while the combined approach yields the highest spBLEU (19.88), the chrF++ score (35.21) is slightly lower than the sentence-only configuration (35.28). This suggests a nuanced trade-off in metric sensitivity: the lexicon provides high-precision constraints that improve exact subword-level overlap (boosting spBLEU), which is otherwise hindered by the 25.9% OOV rate. Conversely, the character-level overlap (chrF++) appears to reach saturation with sentence-level retrieval alone, nearly matching the in-domain reference of 36.17. The slight drop in chrF++ when adding the full lexicon may indicate that the increased context volume introduces minor syntactic noise that outweighs marginal gains in character-level accuracy. Nevertheless, the combined model remains the most robust overall system for recovering performance across both subword and character granularities.
@@ 5.5 Recommendations
Standard practices in low-resource NMT often involve randomly splitting available corpora (e.g., the Bible) into training and test sets (Vázquez et al., 2021; Marashian et al., 2025). While many papers assume that the Bible belongs to a single, religious domain, our analysis shows a marked domain shift within this text, demonstrating that measuring out-of-domain performance is possible even when only the Bible is available as parallel corpus. Therefore, we recommend that low-resource researchers take this into consideration instead of using all verses of the Bible for both train and test, opting instead for document-level holdouts (e.g., distinct books or Testaments) to avoid inflated performance estimates (Khiu et al., 2024). This aligns with recent findings from the WMT 2025 General Translation task Kocmi et al. (2025), which argue that evaluating on "easy" in-domain data masks model brittleness and that robust assessment requires testing on challenging, document-level out-of-domain holdouts.
@@ 6 Conclusion
Crucially, we find that context volume, rather than retrieval algorithm, is the primary driver of performance. We observe that distinct retrieval strategies (lexical vs. semantic) converge to comparable quality levels when normalized for volume. By validating these trade-offs on a language with no digital footprint, we provide a scalable blueprint for accelerating the translation of the Old Testament for thousands of low-resource languages worldwide.
@@ 7 Limitations
This work utilizes data from the Dhao Alkitab (copyright ©2012 Unit Bahasa dan Budaya) and A Grammar of Dhao (Balukh, 2020). The Bible translation is licensed under Creative Commons Attribution-NoDerivatives 4.0 International (CC BY-ND 4.0), which permits redistribution for research purposes provided the text remains unaltered. The grammar book is an Open Access publication (LOT Dissertation Series 570). Our use of these materials for non-commercial linguistic analysis and machine translation evaluation is consistent with these licenses and established Fair Use protocols for academic research. We provide full attribution to the original authors and rights holders in our citations.
@@
Acknowledgements
This research was supported by the Commonwealth through an Australian Government Research Training Program Scholarship (https://doi.org/10.82133/C42F-K220). This research was undertaken using the LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200. Lau was supported by the Australian Research Council under Grant LP210200917. The authors also wish to thank Jermy Balukh for the use of his work, A Grammar of Dhao, which served as a foundational resource for our supplementary data extraction.